OpenStudy (anonymous):

integral of sin^4 (6x) dx

OpenStudy (anonymous):

Take the integral: integral sin^4(6 x) dx For the integrand sin^4(6 x), substitute u = 6 x and du = 6 dx: = 1/6 integral sin^4(u) du Use the reduction formula, integral sin^m(u) du = -(cos(u) sin^(m-1)(u))/m + (m-1)/m integral sin^(m-2)(u) du, where m = 4: = 1/8 integral sin^2(u) du-1/24 sin^3(u) cos(u) Write sin^2(u) as 1/2-1/2 cos(2 u): = 1/8 integral (1/2-1/2 cos(2 u)) du-1/24 sin^3(u) cos(u) Integrate the sum term by term and factor out constants: = -1/24 sin^3(u) cos(u)+1/8 integral 1/2 du-1/16 integral cos(2 u) du For the integrand cos(2 u), substitute s = 2 u and ds = 2 du: = -1/32 integral cos(s) ds-1/24 sin^3(u) cos(u)+1/8 integral 1/2 du The integral of cos(s) is sin(s): = -(sin(s))/32+1/8 integral 1/2 du-1/24 sin^3(u) cos(u) The integral of 1/2 is u/2: = -(sin(s))/32+u/16-1/24 sin^3(u) cos(u)+constant Substitute back for s = 2 u: = u/16-1/32 sin(2 u)-1/24 sin^3(u) cos(u)+constant Substitute back for u = 6 x: = (3 x)/8-1/32 sin(12 x)-1/24 sin^3(6 x) cos(6 x)+constant Which is equal to: Answer: | | = 1/192 (72 x-8 sin(12 x)+sin(24 x))+constant

OpenStudy (anonymous):

As an alternative, you can rewrite the intregrand with the half-angle identity: \[\sin^2\theta=\frac{1}{2}\left(1-\cos(2\theta)\right)\] So, you can rewrite the integrand: \[\begin{align*}\sin^4(6x)&=\left[\frac{1}{2}\left(1-\cos^2(12x)\right)\right]^2\\ &=\frac{1}{4}\left(1-\cos(12x)\right)^2\\ &=\frac{1}{4}\left(1-2\cos(12x)+\cos^2(12x)\right) \end{align*}\] And using the identity involving cosine: \[\cos^2\theta=\frac{1}{2}(1+\cos(2\theta))\] So, \[\begin{align*}\sin^4(6x)&=\frac{1}{4}\left(1-2\cos(12x)+\frac{1}{2}\left(1+\cos(24x)\right)\right)\\ &=\frac{1}{4}\left(\frac{3}{2}-2\cos(12x)+\frac{1}{2}\cos(24x)\right)\\ &=\frac{3}{8}-\frac{1}{2}\cos(12x)+\frac{1}{8}\cos(24x) \end{align*}\] Then term-by-term integration can be easily done.

OpenStudy (anonymous):

I'm sorry i should have told you that my professor does not want us to use the formula, this is what i saw on wolf but I'm trying to integrate it out all without using the formula... so far I'm down to 1/24[u-(2sin(2u)/2)] + integral cos^2 (2u) du

OpenStudy (anonymous):

Does he just want you to use your calculator?

OpenStudy (anonymous):

no no no, this is all done analytically - @SithsAndGiggles this is partially of what i have started to do, I started out using u-sub for 6x then used half-angle identity.

OpenStudy (anonymous):

@sjerman1, did you check my solution, I don't think I use half-angle.

OpenStudy (anonymous):

\[\int\limits \sin^4 (6x) dx \] u=6x du=6dx -> 1/6du=dx \[1/6 \int\limits \sin^4 (u) du\] \[\sin^2 u = \frac{ 1-\cos(2u) }{ 2 } \therefore \sin^4 u = [\frac{ (1-\cos(2u) }{ 2 }]^2 = \frac{ 1-2\cos(2u)+\cos^2 (2u) }{ 4 }\] \[\int\limits\limits \sin^4 (6x) dx = \frac{ 1 }{ 6 } \int\limits \frac{ 1 }{ 4 } (1-2\cos(2u)+\cos^2(2u)) du\] \[\frac{ 1 }{ 6 }\times \frac{ 1 }{ 4 } = \frac{ 1 }{ 24 }[u-\frac{ 2\sin(2u) }{ 2 }] + \int\limits \cos^2 (2u) du\]

OpenStudy (anonymous):

@Zelda I noticed that you didnt use the half angle formula, but my professor does not want me using the formulas such as the reduction formula that you pulled from wolframalpha

OpenStudy (anonymous):

Fair enough. But gotta love wolfram alpha. I hate integrating myself >.>

OpenStudy (anonymous):

Which formulas are you allowed and not allowed to use?

OpenStudy (anonymous):

trig formula, half-angle formula and trig identities. nothing that is like a plug and play formula such as the reduction formula

OpenStudy (anonymous):

\[16×14=124[u−2\sin(2u)2]+∫\cos2(2u)du\] \[\int\limits \cos^2 (2u) du = \int\limits \frac{ 1 }{ 2 } + \frac{ 1 }{ 2 }\cos4u du\] \[\frac{ 1 }{ 24 }[\frac{ 3 }{ 2 }u - \frac{ 2\sin(2u) }{ 2 } +\frac{ 1 }{ 2 } \frac{ \sin(4u) }{ 4 }+c]\] replace 6x for u \[\frac{ 1 }{ 24 }[9x-\frac{ 2\sin(12x) }{ 2 }+\frac{ 1 }{ 8 }\sin(24x)+c]\] there we go, now we are done...